The sum of two integers is 30. Two times the larger integer is 10 less than five times the smaller integer. Find the integers.
Answer by ewatrrr(24785)


Hi
The sum of two integers is 30
Let x and (30-x) represent the smaller and larger number respectively
Question states***
2(30-x) = 5x -10
solving for x
60 - 2x = 5x-10
70 = 7x
x = 10, the smaller number. The larger number is 20. (30-10)
CHECKING our Answer***
40 = 50 - 10